# Short-Form Video Recommendations with Multimodal Embeddings: Addressing Cold-Start and Bias Challenges

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systems
@@ 2. Background
Once interaction data becomes available, personalization is typically introduced using a conventional two-tower architecture (Yang et al., 2020; Covington et al., 2016; Dzhoha et al., 2025), which has become standard for scalable candidate generation. These models learn separate user and item embeddings for efficient retrieval. Given the limitations of early video interaction data, a pragmatic approach is to use non-trainable user embeddings from existing platform models, while training the video tower from scratch on video interactions. Relevance is often defined by videos achieving watch times above a 50% threshold, following industry standards; however, this introduces a bias toward shorter videos (Zhan et al., 2022; Zheng et al., 2022). Combined with position bias and feedback loops, the resulting data can be challenging, especially in immersive feed experiences where position bias is amplified (Liu et al., 2023). Training typically uses dot-product similarity with sigmoid activation, optimized via binary cross-entropy loss, and evaluated using AUC and NDCG metrics.
@@ 4.4. LVLM-as-a-judge evaluation
[…]
@@ 5. Conclusion
Immersive short-form video feeds introduce unique challenges for recommender systems, particularly due to strong position and duration biases that can distort relevance and fairness – especially in new product launches with limited interaction data. Our work demonstrates that conventional supervised approaches, while effective in mature settings, are highly susceptible to these biases and may not generalize well. By leveraging a scalable multimodal retrieval system based on vision-language models and guided by LLMs for evaluation, we effectively addressed these challenges. Our approach enabled robust personalization by mapping user history and video content into a shared semantic space, reducing the impact of position and duration biases. Online experiments confirmed substantial improvements in both content relevance and user engagement, with significant gains in watch-time completion rates and a more balanced distribution of video popularity and duration. These results highlight the importance of foundation models and external evaluation frameworks, such as LVLM-as-a-judge, for building fair and effective recommender systems in immersive video environments.

### v2

@@ Abstract.
copyright: nonecopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Workshop on Evaluating and Applying Recommender Systems with Large Language Models in conjunction with RecSys ’25; September, 2025; Prague, Czech Republicccs: Information systems Recommender systems
@@ 2. Background
Once interaction data becomes available, personalization is typically introduced using a conventional two-tower architecture (Yang et al., 2020; Covington et al., 2016; Dzhoha et al., 2025), which has become standard for scalable candidate generation. These models learn separate user and item embeddings for efficient retrieval. More advanced user models could represent interaction history as a graph of fashion interests, leveraging Siamese graph neural networks (Krivosheev et al., 2021). Given the limitations of early video interaction data, a pragmatic approach is to use non-trainable user embeddings from existing platform models, while training the video tower from scratch on video interactions. Relevance is often defined by videos achieving watch times above a 50% threshold, following industry standards; however, this introduces a bias toward shorter videos (Zhan et al., 2022; Zheng et al., 2022). Combined with position bias and feedback loops, the resulting data can be challenging, especially in immersive feed experiences where position bias is amplified (Liu et al., 2023). Training typically uses dot-product similarity with sigmoid activation, optimized via binary cross-entropy loss, and evaluated using AUC and NDCG metrics.
@@ 4.4. LVLM-as-a-judge evaluation
To complement standard metrics, we used LVLM-as-a-judge (Large Vision-Language Model) for offline evaluation. Qwen 2.5-VL served as an external judge, rating the relevance of top-$k$ ($k=5,10$) recommended videos based on a user’s 12 most recent items of interest. Ratings were assigned on a 5-point scale, from 5 (extremely relevant) to 1 (no relevance). This approach provides an external perspective, especially valuable when user behavior data is limited. The prompt used for LVLM-as-a-judge is summarized in Figure 2.
@@ 5. Conclusion
Immersive short-form video feeds introduce unique challenges for recommender systems, particularly due to strong position and duration biases that can distort relevance and fairness – especially in new product launches with limited interaction data. Our work demonstrates that conventional supervised approaches, while effective in mature settings, are susceptible to these biases and may not generalize well. By leveraging a scalable multimodal retrieval system based on vision-language models and guided by LLMs for evaluation, we effectively addressed these challenges. These results highlight the importance of foundation models and external evaluation frameworks, such as LVLM-as-a-judge, for building fair and effective recommender systems in immersive video environments. Future work will further explore advanced LVLM-based representations to deepen user-video understanding, quantify the reliability of LLM-based judgments (Bhattacharjya et al., 2025), and continue mitigating bias in evolving recommendation scenarios.